# How Network Topology Affects the Strength of Dangerous Power Grid Perturbations

### v2

@@ 2.5 Network motifs
J. Nitzbon et al. [22] have identified some important network motifs relevant to studying power grid stability. These network motifs are described below.
Consider a graph $G=(V,E)$ with vertices $V$ and edges $E$. $G=(V,E)$ is a tree if it is connected and has no cycles. $T^{\prime}=(V^{\prime},E^{\prime})$ is a tree-shaped part of $G$ if it is an induced subgraph of $G$, is a tree and is maximal with the property that there is exactly one node $r\in V^{\prime}$ that has at least one neighbour in $G-T^{\prime}$. The node $r$ of $T^{\prime}$ is called a root node ($R$). The union of all tree-shaped parts of the graph $G$ is called the forest $T$ of the graph $G$. For a tree-shaped part, the depth of node $x$ is the length of the shortest path from node $x$ to the root node.
@@
Nodes that do not belong to the forest part $T$ are called bulk nodes ($B$). Nodes that belong to the forest part and are not root nodes are called non-root nodes. Non-root nodes that have a degree greater than one are called inner tree nodes ($I$). Non-root nodes that have a degree of one are called leaves. Leaves that have a depth of more than one are proper leaves ($P$), and leaves that have a depth of one are called sprouts. Sprouts with an average neighbour degree of less than six are called sparse sprouts ($S_{s}$), and sprouts with an average neighbour degree of more than five are called dense sprouts ($S_{d}$). Fig. 1 shows a network with nodes classified as described above.
We further divide the inner leave nodes into two classes. Class 1 inner tree nodes are inner tree nodes with a degree of two, and class 2 inner leaf nodes are inner tree nodes with a degree greater than two. We also divide dense sprouts into two classes. Class 1 dense sprouts are dense sprouts with a degree of six and seven, and class 2 dense sprouts have a degree greater than six.
@@ 3 Results
We use a random network generator model proposed by Schultz et al. [32] to generate realistic power grid networks to study the effect of network topology on power grid stability. The parameters chosen for this model are $N_{0}=1$, $p=1/5$, $q=3/10$, $s=1/3$, $t=1/10$. $50$ such networks, each consisting of $100$ nodes, were generated. In each network, half of the nodes were taken to be generators, and half of the nodes were taken to be consumers. Grid networks were modelled using equations (1ba) and (1bb), with the following parameters: $K=6$ for every transmission line, $\alpha=0.1$ for every node, $P=1$ for every generator, and $P=-1$ for every consumer. One unit of time in the differential equations corresponds to $0.25$ s. Using this ensemble of grid networks, we investigate the single-node basin stability bound and the single-node survivability bound for every node of all the networks.
Deviations in the grid frequency of power grids are generally kept within ±$\pm 1$ Hz. A deviation of ±$\pm 1$ Hz in the frequency corresponds to a $\omega\approx\pm 1.57$ in the units we have used. We choose two values for the desirable region’s frequency bound $\omega^{+}$. We choose $\omega^{+}=1$ corresponding to an allowed frequency deviation of ±$\pm 0.64$ Hz. Another value $\omega^{+}=5$, corresponding to an allowed frequency deviation of ±$\pm 3.2$ Hz, is chosen to investigate how the survivability bound changes when large frequency deviations are allowed. The region $X^{sn}=[-\pi,\pi]$ represents single-node perturbations in the phase angle. The single-node basin stability bound and single-node survivability bound of node $i$ in a network are computed in the region $X_{i}^{0}$ given by equation (1bg) for single-node perturbations in the region $X^{sn}$. Basin stability bound and survivability bound are computed using tolerances $t_{\beta}=0.95$ and $t_{\sigma}=0.95$, respectively, and the maximum perturbation distance limit is set to $d_{\rm max}=50$. For all basin stability and survivability computations, $n=200$ initial conditions were used, and every initial condition was evolved for time $t=100$. Numerically, a perturbation is counted to be in the basin of attraction if $|\omega_{i}|<1\>\forall i$ at $t=100$.
@@
Fig. 2 shows the relationship between degree and single-node stability. Fig. 2(a) does not show a correlation between basin stability bound and a node’s degree. Nodes with low basin stability bound are not captured by this plot as their values lie below the $15.9\%$ percentile range; these nodes are better examined by looking at the basin stability bound values for different network motifs. Fig. 2(b) shows the dependence of survivability bound on degree — a strong negative correlation is seen. Fig. 2(c) also shows a negative correlation with survivability bound and degree. However, nodes with a degree of one show variable values of survivability bound. Nodes of degree one can be further bifurcated into proper leaves, sparse sprouts, class 1 dense sprouts and class 2 dense sprouts. These node classes are examined to identify vulnerable network structures better.
@@
Fig. 3 shows the average single-node stability values for the different classes of nodes. In Fig. 3(a), we observe that class 1 and class 2 dense sprouts are the least stable nodes out of all the nodes with a degree of one. Additionally, Fig. 3(a) shows that class 1 and class 2 inner tree nodes have low stability bound values. In Fig. 3(b), we observe that root nodes and bulk nodes have the lowest survivability bound values. These nodes are in the interior of networks and are thus generally well-connected and have a higher degree than other nodes. This agrees with the fact that survivability bound is negatively correlated with a node’s degree (Fig. 2(b)). In Fig 3(c), it is observed that dense sprouts have the lowest survivability bound values. Thus, as the allowed frequency deviations increase, dense sprouts move from having relatively high survivability values compared to other nodes to having the lowest survivability values.
@@ 4 Conclusion
We have studied the effect of network topology on single-node stability using these distance-based measures of stability, departing from commonly used volume-based stability quantifiers. Through the study of the single-node stability of an ensemble of synthetic power grids, we have identified vulnerable local network properties of power grids. We have found that inner leave nodes and dense sprouts have the lowest values of basin stability bound. Hence, large perturbations to these nodes should be avoided at all costs. On the other hand, nodes with a high degree have low survivability bound values (when the tolerated frequency deviations in the grid are small (±$\pm 0.64$ Hz)). This means that perturbations to these nodes have the highest chance of resulting in undesirable transients in the grid but may rarely ever result in total loss of synchrony due to such nodes having high basin stability bound values. However, when the tolerated frequency deviations are larger ((±$\pm 3.2$ Hz), dense sprouts have the lowest survivability values. Dense sprouts have the lowest basin stability bound and survivability bound values in this case.
We believe the methods used and the results obtained in this study can prove useful to future research on the stability of power grids and other complex networks.
We thank Dr. Frank Hellmann from Potsdam Institute for Climate Impact Research, for his helpful suggestions. S B acknowledges the J C Bose National Fellowship provided by SERB, Government of India, Grant No. JBR/2020/000049.

### v3

@@ 2.5 Network motifs
J. Nitzbon et al. Nitzbon et al., [2017] have identified some important network motifs relevant to studying power grid stability. These network motifs are described below.
Consider a graph $G=(V,E)$ with vertices $V$ and edges $E$. $G=(V,E)$ is a tree if it is connected and has no cycles. $T^{\prime}=(V^{\prime},E^{\prime})$ is a tree-shaped part of $G$ if it is an induced subgraph of $G$, is a tree and is maximal with the property that there is exactly one node $r\in V^{\prime}$ that has at least one neighbor in $G-T^{\prime}$. The node $r$ of $T^{\prime}$ is called a root node ($R$). The union of all tree-shaped parts of the graph $G$ is called the forest $T$ of the graph $G$. For a tree-shaped part, the depth of node $x$ is the length of the shortest path from node $x$ to the root node.
@@
Nodes that do not belong to the forest part $T$ are called bulk nodes ($B$). Nodes that belong to the forest part and are not root nodes are called non-root nodes. Non-root nodes that have a degree greater than one are called inner tree nodes ($I$). Non-root nodes that have a degree of one are called leaves. Leaves that have a depth of more than one are proper leaves ($P$), and leaves that have a depth of one are called sprouts. Sprouts with an average neighbor degree of less than six are called sparse sprouts ($S_{s}$), and sprouts with an average neighbour degree of more than five are called dense sprouts ($S_{d}$). Fig. 1 shows a network with nodes classified as described above.
In addition to the above classification scheme, we further divide the inner leave nodes into two classes. Class 1 inner tree nodes are inner tree nodes with a degree of two, and class 2 inner leaf nodes are inner tree nodes with a degree greater than two. We also divide dense sprouts into two classes. Class 1 dense sprouts are dense sprouts with a degree of six and seven, and class 2 dense sprouts have a degree greater than six.
@@ 3 Results
We use a random network generator model proposed by Schultz et al. Schultz et al., 2014b to generate realistic power grid networks to study the effect of network topology on power grid stability. The parameters chosen for this model are $N_{0}=1$, $p=1/5$, $q=3/10$, $s=1/3$, $t=1/10$. 5050 such networks, each consisting of $100$ nodes, were generated. In each network, half of the nodes were taken to be generators, and half of the nodes were taken to be consumers. Grid networks were modelled using equations (3) and (4), with the following parameters: $K=6$ for every transmission line, $\alpha=0.1$ for every node, $P=1$ for every generator, and P=1P=-1 for every consumer. One unit of time in the differential equations corresponds to $0.25$ s. Using this ensemble of grid networks, we investigate the single-node basin stability bound and the single-node survivability bound for every node of all the networks.
Deviations in the grid frequency of power grids are generally kept within $\pm 1$ Hz. A deviation of $\pm 1$ Hz in the frequency corresponds to a $\omega\approx\pm 1.57$ in the units we have used. We choose two values for the desirable region’s frequency bound $\omega^{+}$. We choose $\omega^{+}=1$ corresponding to an allowed frequency deviation of $\pm 0.64$ Hz. Another value $\omega^{+}=5$, corresponding to an allowed frequency deviation of $\pm 3.2$ Hz, is chosen to investigate how the survivability bound changes when large frequency deviations are allowed. The region $X^{sn}=[-\pi,\pi]$ represents single-node perturbations in the phase angle. The single-node basin stability bound and single-node survivability bound of node $i$ in a network are computed in the region $X_{i}^{0}$ given by equation (9) for single-node perturbations in the region $X^{sn}$. Basin stability bound and survivability bound are computed using tolerances $t_{\beta}=0.95$ and $t_{\sigma}=0.95$, respectively, and the maximum perturbation distance limit is set to $d_{\rm max}=50$. For all basin stability and survivability computations, $n=200$ initial conditions were used, and every initial condition was evolved for time $t=100$. Numerically, a perturbation is counted to be in the basin of attraction if $|\omega_{i}|<1\>\forall i$ at $t=100$.
@@
Fig. 2 shows the relationship between degree and single-node stability. Fig. 2(a) does not show a correlation between basin stability bound and a node’s degree. Nodes with low basin stability bound are not captured by this plot as their values lie below the $15.9\%$ percentile range; these nodes are better examined by looking at the basin stability bound values for different network motifs. Fig. 2(b) shows the dependence of survivability bound $(\omega^{+}=1)$ on degree — a strong negative correlation is seen. For small perturbations, it is numerically observed that, after the initial phase and frequency disturbances in the systems, oscillations quickly die out. Thus, for a smaller desirable region, as in the case of $\omega^{+}=1$, where perturbations inside the survivability bound are small, the initial frequency fluctuations in the system are responsible for pushing the system out of the region of survival. Consider a single-node perturbation given by $x_{i}^{p}=(\theta_{1}^{s},\>\theta_{2}^{s},\>...,\>\theta_{i}^{p},\>...,\>\theta_{N}^{s},\>0,\>0,\>...,\>0,\>...,\>0)$. Assuming that the time period $\Delta t$ and the frequency deviations $\Delta\omega_{i}$ after the perturbation are sufficiently small, then $\Delta\omega_{i}$ can be approximated as $\Delta\omega_{i}=\dot{\omega}_{i}\bigl{|}_{x=x_{i}^{p}}\,\Delta t$ right after the perturbation $x_{i}^{p}$, before the effect of nonlinearities and damping of the oscillations. In time $\Delta t$, the maximum frequency deviation in the grid can be written as $\max_{i}\{\Delta\omega_{i}\}=\max_{i}\{\dot{\omega}_{i}\bigl{|}_{x=x_{i}^{p}}\}\,\Delta t$. For small single-node perturbations, the expression $\max_{i}\{\dot{\omega}_{i}\bigl{|}_{x=x_{i}^{p}}\}$, where $\dot{\omega}_{i}$ is given by Eq. 4, is positively correlated with a nodes degree, thus making the initial grid frequency deviation $\max_{i}\{\Delta\omega_{i}\}$ higher for nodes with higher degrees. Thus, when $\omega^{+}$ is small, smaller perturbations are required at high-degree nodes to keep frequency deviations under the bound $\omega^{+}$ — this explains the trend observed in Fig. 2(b).
@@
Fig. 2(c) also shows a negative correlation with survivability bound $(\omega^{+}=5)$ and degree. However, nodes with a degree of one take a wide range of survivability bound values. Nodes of degree one can be further bifurcated into proper leaves, sparse sprouts, class 1 dense sprouts, and class 2 dense sprouts. These node classes are examined in Fig. 3.
Fig. 3 shows the average single-node stability values for the different classes of nodes described in section 2.5. In Fig. 3(a), we observe that class 1 and class 2 dense sprouts have the lowest basin stability bound values out of all the nodes with a degree of one. Additionally, Fig. 3(a) shows that class 1 and class 2 inner tree nodes have low basin stability bound values. In Fig. 3(b), we observe that root nodes and bulk nodes have the lowest survivability bound values (with $w^{+}=1$). These nodes are in the interior of networks and thus have a higher degree than other nodes. This agrees with the fact that survivability bound is negatively correlated with a node’s degree (Fig. 2(b)). In Fig 3(c), it is observed that dense sprouts have the lowest survivability bound values (with $w^{+}=5$). Thus, as the allowed frequency deviations increase, dense sprouts move from having relatively high survivability values compared to other nodes to having the lowest survivability values.
The plots in Fig. 3 capture the average single-node stability for different node classes; however, these plots do not account for outlier values that lie below the $15.9\%$ percentile. To investigate the most vulnerable nodes in a network, examining the stability of outlier nodes is required. To examine the stability of the least stable nodes for each node class, we define the following metric known as tail expectation:
$T_{c,\>p}=\mathbb{E}[X_{c}|X_{c}\leq q_{c,\>p}]$ (22)
Here, $X_{c}$ is a random variable that takes on stability values that follow the distribution of the computed stability values of the node class $c$, and $q_{c,\>p}$ is the value of the $p^{\text{th}}$ percentile of the distribution. For the analysis, we choose $p=3\%$, implying $T_{c,\>p}$ is the expected value of stability for the lowest $3\%$ of stability values.
Using this metric, we obtain plots (Fig. 4) for basin stability bound and survivability bound $(\omega^{+}=5)$.
Figure 4: Bar graph showing the tail expectation of single-node stability for bulk nodes ($B$), root nodes ($R$), class 1 inner tree nodes ($I_{1}$), class 2 inner tree nodes ($I_{2}$), proper leaves ($P$), sparse sprouts ($S_{s})$, class 1 dense sprouts ($S_{d1}$) and class 2 dense sprouts ($S_{d2}$). (a) Basin stability bound for each node class. (b) Survivability bound ($w^{+}=5$) for each node class. The black line indicates the confidence interval corresponding to one standard error.
In Fig. 4(a) and Fig. 4(b), the tail expectation of both basin stability bound and survivability bound is the lowest for proper leaves and dense sprouts, indicating that the nodes with the lowest stability are proper leaves and dense sprouts.
In Fig. 5, we have perturbed the most unstable proper leave nodes and dense sprouts with perturbations of strengths equal to their tail expectations. It is observed that the perturbation to the proper leave node leaves the proper leave oscillating with a distinctively different frequency, while the other nodes in the network oscillate closely around the grid frequency. This implies that, after a perturbation, control to the proper leave alone could lead to the grid being resynchronized. On the other hand, the perturbation to the dense sprout causes multiple nodes to oscillate at different frequencies leaving the grid in a desynchronized state described by J Nitzbon et al. Nitzbon et al., [2017].
Figure 5: Plot showing the frequency ($\omega$) versus time ($t$) of all nodes in a network for a perturbation to a (a) proper leave and a (b) dense sprout. The blue curve is the trajectory of the perturbated node, while the grey curves are the trajectories of the other nodes in the network.
Previous studies that have examined power grid stability using a Kuramoto-like model have identified the heightened vulnerability of inner leaves (based on betweenness centrality) Menck et al., [2014], as well as dense sprouts Nitzbon et al., [2017]. However, the vulnerability of proper leaves has not been observed before. In this study, using probabilistic distance-based stability quantifiers, we have identified proper leaves and dense sprouts to be the most vulnerable in the case of large perturbations and large allowed grid frequency deviations.
@@ 4 Conclusion
We have studied the effect of network topology on single-node stability using these distance-based measures of stability, departing from commonly used volume-based stability quantifiers. Through the study of the single-node stability of an ensemble of synthetic power grids, we have identified vulnerable local network properties of power grids. We have found that proper leaves and dense sprouts have the lowest values of basin stability bound. Hence, large perturbations to these nodes should be avoided at all costs. On the other hand, nodes with a high degree have low survivability bound values (when the tolerated frequency deviations in the grid are small ($\pm 0.64$ Hz)). This means that perturbations to these nodes have the highest chance of resulting in undesirable transients in the grid but may rarely ever result in total loss of synchrony due to such nodes having high basin stability bound values. However, when the tolerated frequency deviations are larger (($\pm 3.2$ Hz), proper leaves and dense sprouts have the lowest survivability values.
In this paper, we have used distance-based nonlinear stability methods on power grids, however, they are readily applicable to other complex-network dynamical systems. We believe that the methods used and the results obtained in this study can be useful in future research on the stability of power grids and other complex networks.
Acknowledgments
We thank Dr. Frank Hellmann from the Potsdam Institute for Climate Impact Research for his helpful suggestions. S B acknowledges the J C Bose National Fellowship provided by SERB, Government of India, Grant No. JBR/2020/000049.
ORCID